Find the slope of the line passing through the points (0, -4) and (-10, 9).
a) 10/13
b) -10/13
c) 13/10
d) -13/10

Answer by rfadrogane(214)

Find the slope of the line passing through the points (0, -4) and (-10, 9).
a) 10/13
b) -10/13
c) 13/10
d) -13/10

Sol'n:

m = (y2-y1)/(x2-x1)
m = (9+4)/(-10-0)
m = -13/10----answer